OpenStudy (anonymous):

11 *radical 6 divided by radical 98

OpenStudy (anonymous):

do you mean? \[\frac{ 11*\sqrt{6} }{ \sqrt{98} }\]

OpenStudy (anonymous):

if yes then what you do is you need to first simplify the radical \[\frac{ 11*\sqrt{6} }{ \sqrt{98} }*\frac{ \sqrt{98} }{ \sqrt{98} }\]

OpenStudy (anonymous):

\[\sqrt{98}*11=77\sqrt{2}\] \[\sqrt{98}*\sqrt{6}=14\sqrt{3}\]

OpenStudy (anonymous):

so that is your top now just multiply those two together and then find your bottom

OpenStudy (anonymous):

yes thats what i meant

OpenStudy (anonymous):

thanks a million